# A Reliable Framework for Human-in-the-Loop Anomaly Detection in Time Series

### v2

@@ II-C1 Time-series Context Similarity Calculation
II-C2 Attribution Similarity Calculation
Given a fully convolutional network (FCN), $M$, trained for time series binary classification task (i.e., normal or abnormal), an input time series sequence, $T_{Q}$, and the predicted class of the $M$ for input $T_{Q}$, $c(T_{Q})$, the attention score of $T_{Q}$ is calculated as
$a(T_{Q})=CAM(M,T_{Q},c(T_{Q})),$ (2)
where $a(T_{Q})$ is a vector with the same dimension as $T_{Q}$, and each position in $a(T_{Q})$ corresponds to the degree of dependence of model $M$ on the corresponding dimension and time point when it makes the prediction $c(T_{Q})$. Thus $a(T_{Q})$ represents the model attribution. Please refer [21] to the specific calculation process of CAM. For input sequences $T_{Q}$ and $T_{P}$, the attribution similarity, $d_{cos}(a(T_{Q}),a(T_{P}))$, is computed using the cosine distance as:
@@ II-C2 Attribution Similarity Calculation
where $\|a(T_{Q})\|_{2}$ and $\|a(T_{P})\|_{2}$ denote the L2 norm of vectors $a(T_{Q})$ and $T_{a}(P)$, respectively. Consistent with $d_{DTW}$, a smaller cosine distance indicates a higher similarity.
II-C3 Attribution-Aware Clustering and Summarization
To aggregate data with similar context and also similar model attribution together, we construct the following distance $m_{QP}$ to measure the similarity between two instances $T_{Q}$ and $T_{P}$:
@@ II-C3 Attribution-Aware Clustering and Summarization
where $\alpha$ is the weight to balance the data context similarity $d_{DTW}(T_{Q},T_{P})$ and the model attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$. In this work, we set $\alpha=0.5$ in all case studies. Thus, for all of the $N$ time series instances, we can get the following distance matrix:
$M_{dist}=\begin{bmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N,1}&m_{N,2}&\cdots&m_{N,N}\end{bmatrix}.$ (5)
Finally, we use UMAP [27] to plot $M_{dist}$ and apply K-means [28] to generate $K$ clusters:
@@
where $C=\{c_{1},c_{2},\dots,c_{K}\}$ is the set of clusters, UMAP$(M_{dist})$ is the 2D representation of the data and $K$ is the desired number of clusters. As a follow-up step, we visually summarize the resulting clusters as an integrated visualization for each cluster, which presents the common patterns of time sequences and model attributions in the current cluster, empowering users to grasp model behavior patterns more efficiently.
II-D Annotation and Propagation
Building upon the aforementioned method, at the “Annotation and Propagation” phase, we To effectively evaluate such spurious attention, we introduce a metric termed “Spuriousness score”, $S[c_{i}]$, ranging from 0 to 1, signifying the probability range from the least to the most likelihood of spurious attention occurrence. Furthermore, we deploy the Label Propagation [29], which automatically estimates probabilities of the unannotated clusters’ spuriousness based on users’ annotations and neighboring clusters $N(c_{i})$ in the representation space.
The representation space is generated by UMAP as described in Sec. II-C, where we use the adjacency matrix, $A(i,j)$, to denote the distance between clusters’ feature representations of $c_{i}$ and $c_{j}$. Notably, users only need to provide an annotation for one cluster (Algo. 1, Line 6), designating it as either “correct attention” ($S[c_{i}]=0$) or “spurious attention” ($S[c_{i}]=1$). Our algorithm (Algo. 1, Line 8-15) then propagates this information to compute Spuriousness scores for other clusters. Subsequently, these scores are showcased in the Cluster Information window (Fig. 3  B).
The “Spuriousness score” yields two noteworthy advantages. Primarily, they facilitate a more streamlined exploration process of clusters, with potentially spurious correlations being conspicuously highlighted. This is a pivotal step in aiding users to identify and assess problematic attention. Secondly, following users’ validation, these scores are leveraged to determine the problematic clusters crucial for issue mitigation, as we elucidate in the following section.
@@ II-D Annotation and Propagation
4:
5:$\mathbf{S}$ : Vector of spuriousness scores for each cluster.
6:Initialize $\mathbf{S}$ with -1 for all clusters.
7:for each cluster $c_{i}$ in $C$ with a user label do
8:     $S[c_{i}]$ = label value of $c_{i}$
9:while unlabeled clusters exist in $C$ do
10:     for each unlabeled cluster $c_{i}$ in $C$ do
11:         $\text{WeightedSum}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})\times S[c_{j}]$
12:         $\text{TotalWeight}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})$
13:         $S[c_{i}]=\frac{\text{WeightedSum}(c_{i})}{\text{TotalWeight}(c_{i})}$
14:return $\mathbf{S}$
II-E Model Enhancement
An effective strategy for addressing detected spurious correlations in machine learning (ML) models involves model retraining or finetuning, which reduces the model’s resilience over potential bias while keeping the architecture intact. Retraining the model on a noise-corrupted dataset has demonstrated its efficacy in diminishing the model’s reliance on spurious features [30]. To counteract the influence of spurious features, we harness the masking strategy to shift and correct the model’s attention. The idea is conceptually simple: first, mask the parts of the data that receive spurious attention, then encourage the model to infer how to make decisions with the rest of the data. This idea has been used as a self-supervised pre-training method and spuriousness feature elimination method in Computer Vision (CV) and Natural Language Processing (NLP) fields. In our work, we adjust the Core Risk Minimization (CoRM) methodology [30] in the CV domain into time series data to introduce controlled masks into the annotated spurious attention time points by injecting random Gaussian noise into the data. This is also a fundamental data augmentation strategy called jittering for time series data that can help the model become more robust to variations and be used in masked autoencoder for time series forecasting task [31].
The integration of CoRM into HILAD entails several steps. First, after users annotate the data, we extract a set of instances, $T_{S}$, that display heightened probabilities of spuriousness, indicating undesirable attention. We also extract a set of instances, $T_{C}$, that display heightened probabilities of correctness, indicating correct attention. For the instances in $T_{S}$, pixel attributions (specifically, CAM masks) accentuate the spurious regions. We employ these masks to introduce random Gaussian noise to the identified spurious regions. Mathematically, for an individual time sequence $T_{i}\in T_{S}$, this operation can be succinctly represented as
@@ II-E Model Enhancement
wherein $T_{i}$ is the input time series data, $\mathbf{m}$ denotes the CAM mask, $z$ represents the generated Gaussian noise matrix, and $\odot$ denotes the Hadamard product. Each variable is of the same dimensions as the input time series data. Similarly, for $T_{j}\in T_{C}$, we add random Gaussian noise to the non-highlighted areas to enhance the model’s attention to abnormal patterns:
@@
After substituting the original data with these spuriousness-masked and correctness-enhanced data, we proceed to retrain the model. Subsequently, an evaluation is conducted to determine the extent to which spurious correlations have been alleviated. In Sec. III-D, we elaborate on the evaluation metrics utilized to quantify the effectiveness of our method in mitigating spurious correlations.
III Experimental Design
@@
III-A Task
@@
In the experiment, participants were asked to interact with the interface of HILAD to evaluate a classification model trained on a specific dataset, containing both normal and abnormal time sequences. Specifically, we encouraged them to use as many visual components as possible, follow the system guidance to verify the highlighted issues, interpret them with the provided information, and annotate the confirmed errors in a restricted time. In each experiment, participants were only allowed to perform a single model retraining to better control experiment time and quantify the evaluation results.
@@
With the awareness of the HILAD workflow, participants explored the general model behavior concerning different clusters of data through the Cluster Information window, where they also investigated further details from the Instance Information window that updated simultaneously with their interactions. They were asked to mark and annotate detection issues, such as a model using normal time points to determine an abnormal time series, which could be achieved by simply clicking the “Spurious” button in the Cluster Information window. At this point, all instances corresponding to this cluster were marked as spurious. Participants could further adjust the spuriousness labels of single instances in the Instance Information window. Similarly procedure could be conducted to label and adjust a cluster related to correct attention. As long as one cluster is annotated as spurious or correct, our algorithm will approximate the attention correctness of other clusters according to their similarities in the representation space, resulting in a spuriousness approximation score for every cluster. Participants can click the “spuriousness” button provided in our interface to rank clusters according to this score, allowing for adjustments and confirmation of the approximated model issues. When the participants confirmed the first round of issue inspection could be concluded, they would click the retrain button, and our system performed the model retraining based on the participants’ feedback.
III-B Procedure
In the following, we introduce the experiment’s procedure. The overall practical accomplishment of the experiment took between 30 and 45 minutes.
III-B1 Introduction and Preparations
@@
Before the experiment begins, participants receive detailed instructions explaining the experiment’s purpose and overall structure before it begins. Next, they are introduced to the key concepts involved in time-series classification tasks, including metrics such as confidence and accuracy, possible abnormal patterns, and spurious model attention. Subsequently, an in-depth orientation to HILAD is provided, highlighting its various components and functionalities. This orientation covers the System Menu, Representation Space, Cluster Information, and Instance Information, elucidating their respective roles in detecting anomalies. The primary goal communicated to the participants is to employ HILAD for monitoring effectively and analyzing the model’s behavior, focusing on its attention mechanisms.
III-B2 Familiarization Procedure
@@
III-B3 Experimental Run
In the experimental run, participants were introduced to the HILAD with two space domain-related datasets that mimic its real-world use for anomaly detection. Their task was to critically analyze the behavior of the classification model across the datasets. The main variable in this phase was the dataset itself, and the sequence of dataset presentation varied systematically among participants. This approach aimed to mitigate any potential bias from learning curve effects. We executed three trials for each dataset to further effectively compare the distinctions between the automated algorithm and the human-AI collaboration method. These trials varied in the scope of information provided to the participants: the first trial provided access to only the System Menu window, the second trial included System Menu, Cluster Information, and Representation Space windows, and the third trial offered a comprehensive view with the complete all of the four windows. The first trial in our study exemplifies the application of the automated algorithm, while the latter two trials represent the human-AI collaboration methods. To ensure the integrity of our results and eliminate any potential biases associated with the sequence of exposure, the order of these trials was randomized for each participant. This approach was essential to assess the impact of varying levels of information availability on the performance and decision-making processes in the context of human-AI interaction.
III-B4 Postprocessing
@@
III-C Participants
This experiment was approved by the Institutional Review Board at the University of California, Davis. We involve 15 participants in total, of which 13 were male and 2 were female. All participants are graduate students between the ages of 23 to 30 years old. All participants passed the test to ensure they could accurately identify and annotate the abnormal pattern. All completed the entire experiment on two provided datasets.
III-D Performance Evaluation Metrics
We measure the effectiveness of our proposed framework in terms of classification and attention performance. From the perspective of classification performance, we use accuracy ($Acc$), precision ($P$), recall ($R$), and $F_{1}$ score to evaluate the performance of our method comprehensively. The details of the metrics are listed as follows:
$Acc=\frac{TP+TN}{TP+TN+FP+FN},P=\frac{TP}{TP+FP},$
$R=\frac{TP}{TP+FN},F_{1}=\frac{2\cdot P\cdot R}{P+R},$
where TP, TN, FP, and FN are four classification results, denoting true positive, true negative, false positive, and false negative, respectively. $Acc$ is a measure of the overall correctness of a classification model. $P$ represents the ability of the method to distinguish anomalies, and a higher $P$ accurately indicates fewer false anomalies. $R$ is to determine the ability of the method to detect all the anomalies. Since The $P$ and $R$ metrics often appear contradictory, the $F_{1}$ score is a metric for comprehensively considering these two indicators and providing a balanced overall performance of the method. When the data set has imbalanced classes, F1 Score might be a better metric than accuracy.
In the perspective of attention accuracy performance, we adjust the relevance accuracy ($RA$) proposed in [23] to quantify how accurate a model is at locating ground-truth time points at which the abnormal behavior occurs. The core idea of relevance accuracy is to compare the $k$ most attention time points for a prediction with the ground-truth abnormal time points. For a given time sequence $T_{i}$, let $G(T_{i})$ denote the set of continuous or discontinuous ground-truth abnormal time points, where $|G(T_{i})|$ is the number of ground-truth time points, and $R_{i}(k)$ denote the $k$ time steps that get the most attention for the prediction of the model. To apply this metric to data with varying numbers of ground-truth time points, we set $k=(1+M\%)\times|G(T_{i})|$ and denote $R_{i}(k)$ as $R_{i}(M\%)$ in our work. In this analysis, the sequence of the ground-truth time points is disregarded, given that each ground-truth time point is treated with equal significance. Optimal relevance accuracy is attained when all elements of $G(T_{i})$ are encompassed within $R_{i}(M\%)$, ideally with the minimal possible value of $M$. For a time sequence $T_{i}$, $RA(T_{i})$ can be calculated as:

### v4

@@ 3.2.1 Time series Context Similarity Calculation
3.2.2 Attribution Similarity Calculation
Given a model $\mathcal{M}$, trained for TS binary classification task (i.e., normal or abnormal), an input time sequence, $T_{Q}$, and the predicted class of the $\mathcal{M}$ for input $T_{Q}$, $c(T_{Q})$, the attention score of $T_{Q}$ is calculated as
$a(T_{Q})=CAM(\mathcal{M},T_{Q},c(T_{Q})),$ (2)
where $a(T_{Q})$ is the model attribution mask, which is a one-dimensional vector with the same length as $T_{Q}$. For input sequences $T_{Q}$ and $T_{P}$, the attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$ is computed by:
@@ 3.2.2 Attribution Similarity Calculation
where $\|a(T_{Q})\|_{2}$ and $\|a(T_{P})\|_{2}$ denote the L2 norm of vectors $a(T_{Q})$ and $T_{a}(P)$, respectively. For this cosine distance, similar to $d_{DTW}$, a smaller value indicates a higher similarity.
3.2.3 Attribution-Aware Clustering and Summarization
To aggregate data with both similar context and similar model attribution, we construct the following aggregated distance $m_{QP}$ to measure the similarity between two instances $T_{Q}$ and $T_{P}$:
@@ 3.2.3 Attribution-Aware Clustering and Summarization
where $\alpha$ is the weight to balance the data context similarity $d_{DTW}(T_{Q},T_{P})$ and the model attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$. Both distances are normalized to the range $[0,1]$ to ensure scale compatibility and prevent either term from dominating the aggregated distance. In this work, we set $\alpha=0.5$. Thus, for all of the $N$ TS instances, we can get a distance matrix $M_{dist}\in\mathbb{R}^{N\times N}$.
Finally, we use UMAP McInnes et al. (2018) to project $M_{dist}$ and apply K-means Krishna and Murty (1999) to generate $K$ clusters:
@@
where $C=\{c_{1},c_{2},\dots,c_{K}\}$ is the set of clusters, UMAP$(M_{dist})$ is the 2D representation of the data, and $K$ is the desired number of clusters determined by the elbow method. As a follow-up step, we visually summarize the resulting clusters as an integrated visualization for each cluster, which presents the common patterns of time sequences and model attributions in the current cluster, empowering users to grasp model behavior patterns more efficiently.
3.3 Annotation and Propagation
Building upon the aforementioned method, at the “Annotation and Propagation” phase, we define a metric termed “Spuriousness score", $S[c_{i}]$, to evaluate the presence of spuriousness. This score ranges from 0 to 1, representing the spurious probability from the least to most likely. Furthermore, the Label Propagation algorithm Zhou et al. (2003) is deployed to automatically estimate probabilities of the unannotated clusters’ spuriousness based on users’ annotations and neighboring clusters $N(c_{i})$.
The representation space is generated by UMAP as described in Sec. 3.2, where we use the adjacency matrix, $A(i,j)$, to denote the distance between clusters’ feature representations of $c_{i}$ and $c_{j}$. Notably, users only need to provide binary annotation for a cluster, designating it as either “correct attention" ($S[c_{i}]=0$) or “spurious attention" ($S[c_{i}]=1$). Our algorithm (Algo. 1, Line 6-14) then propagates this information to compute Spuriousness scores for other clusters. Subsequently, these scores are showcased in the Cluster Information window (Fig. 3  B).
@@ 3.3 Annotation and Propagation
4:$\mathbf{Y}$: Initial labels, where $Y[c_{i}]\in\{0,1\}$ for labeled clusters and $-1$ otherwise.
5:
6:$\mathbf{S}$: Vector of propagated spuriousness scores for each cluster.
7:Initialize $\mathbf{S}\leftarrow\mathbf{Y}$
8:for iteration = 1 to max_iter do
9:     $\mathbf{S}^{\text{new}}\leftarrow\mathbf{S}$
10:     for each cluster $c_{i}\in C$ do
11:         if  $Y[c_{i}]\neq-1$  then
12:              continue
13:         $\text{WeightedSum}\leftarrow\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})\times S[c_{j}]$
14:         $\text{TotalWeight}\leftarrow\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})$
15:         $S^{\text{new}}[c_{i}]\leftarrow\frac{\text{WeightedSum}}{\text{TotalWeight}}$
16:     $\mathbf{S}\leftarrow\mathbf{S}^{\text{new}}$
17:return $\mathbf{S}$
The “Spuriousness score" yields two noteworthy advantages. Primarily, they facilitate a more streamlined exploration process of clusters, with potentially spurious correlations being conspicuously highlighted. This is a pivotal step in aiding users to identify and assess problematic attention. Secondly, following users’ validation, these scores are leveraged to determine the problematic clusters crucial for issue mitigation, as we elucidate in the following section.
3.4 Model Enhancement
An effective strategy for addressing detected spurious correlations involves model retraining or finetuning, which reduces the model’s resilience over potential bias while keeping the architecture intact. Specifically, the Core Risk Minimization (CoRM) methodology Singla et al. (2022) that retrains the model on a noise-corrupted dataset has demonstrated its efficacy in spuriousness mitigation for vision models. In our work, we adapt CoRM to TS data, which harnesses the masking strategy by injecting random Gaussian noise into the spurious time points, aiming to shift and correct the model’s wrong attention by encouraging it to infer decisions with the rest of the data. The idea aligns with a data augmentation strategy called jittering for time series data, which can help the model become more robust to variations and be used in masked autoencoders for time series forecasting Tang and Zhang (2022).
The integration of CoRM into HILAD entails several steps. We first extract a set of instances with high spuriousness scores, $T_{S}$, indicating undesirable attention. We also extract a set of instances with the lowest spuriousness, $T_{C}$, indicating correct attention. The attribution masks of $T_{S}$ highlight spurious regions, which are used to occupy such regions with random Gaussian noise accordingly. For an individual time sequence $T_{i}\in T_{S}$, this operation can be mathematically represented as
@@ 3.4 Model Enhancement
where $\mathbf{m}$ denotes the CAM mask, $z$ represents the generated Gaussian noise matrix, and $\odot$ denotes the Hadamard product. Similarly, for $T_{j}\in T_{C}$, we add random Gaussian noise to the non-highlighted areas to enhance the model’s attention to abnormal patterns:
@@
After substituting the original data with these spuriousness-masked and correctness-enhanced data, we proceed to retrain the model. Subsequently, an evaluation is conducted to measure the mitigation of spurious correlations. In Sec. 4.6, we elaborate on the evaluation metrics utilized to quantify the effectiveness of our method in mitigating spurious correlations.
4 Experimental Design
@@
4.1 Datasets
Mars Science Laboratory (MSL) Dataset Hundman et al. (2018) corresponds to the sensor and actuator data for the Mars rover. The dataset provides the ground truth anomaly masks that marks the time point when the abnormal behavior occurs. Since this dataset is known to have many trivial sequences Wu and Keogh (2021), to make a best-faith effort to eliminate mislabeled time points, we build a new dataset based on the two nontrivial ones (A2 and A4). Specifically, we randomly divide the original time series into 1000 equal-length sequences, all of 800 lengths, where the ratio of normal to abnormal data is 3:2.
Simulation Testbed for Exploration Vehicle ECLSS (STEVE) Deng et al. (2023) is a simplified single-bed $CO_{2}$ removal system of the Carbon Dioxide Removal Assembly (CDRA) onboard the International Space Station (ISS), at the University of Colorado Boulder. A corresponding Simulink model is provided to simulate multiple failure modes similar to the STEVE testbed. In our study, we conducted a simulation consisting of four cycles, each comprising 80 minutes of $CO_{2}$ adsorption followed by 80 minutes of desorption. During the third cycle, a simulated leak failure is introduced at a random time point in Valve 1, resulting in observable anomalies in three sensors: Bed Temperature, $CO_{2}$ Concentration, and Flow Rate. These sensor sequences are treated collectively. We generated 1000 sets of sequences with a ratio of 4:1 between normal and abnormal data. Additionally, we marked time points at which abnormal behaviors occurred in the three sensors as ground truth masks.
KPI anomaly detection for AIOps (KPI) dataset Li et al. (2022b) is a real-world, large-scale dataset for evaluating anomaly detection systems in industrial AIOps settings. It contains time series with noisy, high-variance patterns and diverse anomaly types, making it suitable for benchmarking real-world deployment performance. For our experiments, we construct equal-length sequences of 800 time points from four KPI categories with the highest anomaly ratios, resulting in 1,200 time series instances. Additional dataset details are provided in the Appendix.
4.2 Baselines
Baseline models. We involve two baseline models in our experiments. One is a fully convolutional network (FCN) Wang et al. (2017), which has demonstrated effectiveness on time series classification. The other is TST Zerveas et al. (2021), a transformer-based model for multivariate time series anomaly detection. These two models have distinct architectures for evaluating our method with various baselines. Note that HILAD is fully applicable to other time series anomaly classification models with different architectures Boniol et al. (2023); Liu et al. (2021b); Yang et al. (2021). In addition to baseline models, we compare HILAD with the following spuriousness mitigation strategies using data augmentation:
•  Random aug.   Randomly select $\Gamma$ instances from the training set for model fine-tuning.
•  Random mask.   Select $\Gamma$ instances proposed by our system, but mask random time points.
•  Cluster mask.   The $\Gamma$ instances proposed by our system only concede the participant’s operation in the Cluster Information window, which means the instance-level operations are discarded. This method uses the same masking strategy as ours.
4.3 Participants
This experiment was approved by the Institutional Review Board. We involve 15 participants in total, of which 13 were male and 2 were female. All participants are graduate students between the ages of 23 to 30 years old. All participants passed the test to ensure they could accurately identify and annotate the abnormal pattern.
4.4 User Tasks
@@
In the experiment, participants were asked to interact with the interface of HILAD to evaluate a classification model trained on a specific dataset, containing both normal and abnormal time sequences. Specifically, we encouraged them to use as many visual components as possible, follow the system guidance to verify the highlighted issues, interpret them with the provided information, and annotate the confirmed errors within a restricted time. In each experiment, participants were only allowed to perform model retraining once to better quantify the results.
@@
With the awareness of the HILAD workflow, participants explored the general model behavior concerning different clusters of data through the Cluster Information window, where they also investigated further details from the Instance Information window that updated simultaneously with their interactions. They were asked to mark and annotate detection issues, such as a model using normal time points to determine an abnormal time series, which could be achieved by simply clicking the “Spurious" button in the Cluster Information window. At this point, all instances corresponding to this cluster were marked as spurious. Participants could further adjust the spuriousness labels of single instances in the Instance Information window. Similarly procedure could be conducted to label and adjust a cluster related to correct attention. As long as one cluster is annotated as spurious or correct, our algorithm will approximate the attention correctness of other clusters according to their similarities in the representation space, resulting in a spuriousness approximation score for every cluster. Participants can click the “spuriousness" button provided in our interface to rank clusters according to this score, allowing for adjustments and confirmation of the approximated model issues. When the participants confirmed the first round of issue inspection could be concluded, they would click the retrain button, and our system performed the model retraining based on the participants’ feedback.
4.5 Procedure
In the following, we introduce the procedure of our user study.
4.5.1 Introduction and Preparations
@@
Before the experiment begins, participants receive detailed instructions explaining the experiment’s purpose and overall structure before it begins. Next, they are introduced to the key concepts involved in time series classification tasks, including metrics such as confidence and accuracy, possible abnormal patterns, and spurious model attention. Subsequently, an in-depth orientation to HILAD is provided, highlighting its various components and functionalities. This orientation covers the System Menu, Representation Space, Cluster Information, and Instance Information, elucidating their respective roles in detecting anomalies. The primary goal communicated to the participants is to employ HILAD for monitoring effectively and analyzing the model’s behavior, focusing on its attention mechanisms.
4.5.2 Familiarization Procedure
@@
4.5.3 Experimental Run
In the experimental run, participants’ tasks were to critically analyze the behavior of each classification model across three datasets. We executed three trials for each dataset to further effectively compare the distinctions between the automated algorithm and the human-AI collaboration method. These trials varied in the scope of information provided to the participants: the first trial provided access to only the System Menu window, the second trial included System Menu, Cluster Information, and Representation Space windows, and the third trial offered a comprehensive view with the complete all four windows. The first trial in our study exemplifies the application of the automated algorithm, while the latter two trials represent the human-AI collaboration methods. To ensure the integrity of our results and eliminate any potential biases associated with the sequence of exposure, the order of these trials was randomized for each participant. This approach was essential to assess the impact of varying levels of information availability on the performance and decision-making processes in the context of human-AI interaction.
4.5.4 Postprocessing
@@
4.6 Performance Evaluation Metrics
We measure the effectiveness of our proposed framework in terms of classification and attention performance. From the perspective of classification performance, we use the standard Accuracy, Precision, Recall, and $F_{1}$ Score to evaluate the performance of our method comprehensively.
In the perspective of attention accuracy performance, we adjust the relevance accuracy ($RA$) proposed in Wickstrøm et al. (2020) to quantify how accurate a model is at locating ground-truth time points at which the abnormal behavior occurs. The core idea of relevance accuracy is to compare the $k$ most attention time points for a prediction with the ground-truth abnormal time points. For a given time sequence $T_{i}$, let $G(T_{i})$ denote the set of continuous or discontinuous ground-truth abnormal time points, where $|G(T_{i})|$ is the number of ground-truth time points, and $R_{i}(k)$ denote the $k$ time steps that get the most attention for the prediction of the model. To apply this metric to data with varying numbers of ground-truth time points, we set $k=(1+M\%)\times|G(T_{i})|$ and denote $R_{i}(k)$ as $R_{i}(M\%)$ in our work. In this analysis, the sequence of the ground-truth time points is disregarded, given that each ground-truth time point is treated with equal significance. Optimal relevance accuracy is attained when all elements of $G(T_{i})$ are encompassed within $R_{i}(M\%)$, ideally with the minimal possible value of $M$. For a time sequence $T_{i}$, $RA(T_{i})$ can be calculated as: